OpenStudy (anonymous):

express each number in terms of i the suare root of negitive numbers -32 -18 -1/9

OpenStudy (anonymous):

since \(-32=16\times 2\times (-1)\) you have \[\sqrt{32}=\sqrt{16}\sqrt{2}\sqrt{-1}=4\sqrt2 i\]

OpenStudy (anonymous):

\(\sqrt{-18}\) is similar